the position of a particle going along a straight line is x1=100 metres at 10:00 A.M. and x2 = 200 metres at 10:30 P. M. Find the velocity of the particle
1
Expert's answer
2017-10-20T11:45:06-0400
x_1=100 m
x_2=200 m
t_1=10:00 AM
t_2=10:30 PM
The velocity
v=Δx/Δt=(x_2-x_1)/( t_2- t_1 )=(200-100)/12.5=(100 m)/(12.5 hr)=8 m/hr